1.    It is estimated that the Chernobyl disaster released 6.0 Megacurie of 137Cs (half life 30.2yrs) into the environment.

a)   How much activity does Cs have in the year 2021?

b)   Calculate the mass of Cs released?


1
Expert's answer
2021-10-12T12:22:04-0400

We can calculate the mass released using Avogadro’s number and the concept of a mole if we can first find the number of nuclei N released. Since the activity R is given, and the half-life of 137Cs is found to be 30.2 years, we can use the equation

R = 0.693 / Nt1/2 to find N.


Solving the equation R = 0.693 / Nt1/2 for N gives N = Rt1/2 / 0.693


Entering the given values yields

N = ((6.0 MCi) x (30.2 y)) / 0.693


Converting curies to becquerels and years to seconds, we get

N = ((6.0 × 106 Ci) (3.7 × 1010 Bq/Ci) (30.2 y) (3.16×107 s/y)) / 0.693 = 3.1×1026


One mole of a nuclide AX has a mass of A grams, so that one mole of 137Cs has a mass of 137 g. A mole has 6.02 ×1023 nuclei. Thus the mass of 137Cs released was

m = (137 g / 6.02×1023) (3.1×1026) = 70 × 103 g = 70 kg